Design the two primers (6 nucleotides), and add an EcoR1 cut site to the forward primer (the 6-letter sequence), and BamH1 cut site to the reverse primer. Look up the cut site sequence.
Label the restriction enzyme cut sites and add two overhang bases. Cut sites are included in the Tm.

5’-CTATGAGGTCCTGCGTTAGTGTTACC-3’

Forward: Tm:

Reverse: Tm:

Annealing temperature:
Answer #1

what is a primer ?

A Primer s a short strand of RNA or DNA (generally about 18-22 bases) that serves as a starting point for DNA synthesis. It is required for DNA replication because the enzymes that catalyze this process, DNA polymerases, can only add new nucleotides to an existing strand of DNA. The polymerase starts replication at the 3'-end of the primer, and copies the opposite strand

Generally,.primer should have 18-20 nucleotides of overlap with template strand.

5’ primer (forward primer) should align to the first 18-20 nucleotides of the sense strand of the gene ('normal' strand, complement to antisense strand). It should also include the recognition site of the desired restriction enzyme at the 5' end (upstream) of the first nucleotide of overlap

3’ primer (reverse primer) should align to the last 18-20 nucleotides of the anti-sense strand . Make sure a stop codon (TAA (recommended), TAG, TGA) is present if needed. The primer should also include the recognition site of the desired restriction enzyme at the 3' end (downstream) of the last nucleotide of overlap and downstream of the stop codon. Take the reverse-complement of the sequence to make it align to the anti-sense strand.

OUR TEMPLATE: 5’-CTATGAGGTCCTGCGTTAGTGTTACC-3’

FORWARD PRIMER: 5' -CTATGAGGT CCTGCGTTAGTGTTA-3'

REVERSE PRIMER: 5' -GGTAACACTAACGCAGGACCTCAT-3'

FORWARD PRIMER WITH EcoR1: 5' -CTATGAGGTCC TGTTAG TGTTA-3'

REVERSE PRIMER WITH BAMH1: 5' -GGTAACACTAACGCA GGACCT CAT-3'

To calculate the melting temperature of a primer:
Tm (°C) =(G+C)*4+(A+T)*2

Tm of forward primer : Tm=(7+4)*4 +(3+7)*2

=11*4 +10*2

=44+20 =64

Tm of reverse primer : Tm=(5+6)*4 +(7+3)*2

   =11*4+10*2

   =44+20 = 64